One car traveling 40 Km/hr left a certain place 4 hours later than another car traveling in the same direction at the rate of 30 km/hr. In how many hours will the faster car overtake the other?
Answer by checkley77(12569)

D=RT
D=30T
D=40(T-4)
SEEING AS THE 2 DISTANCES ARE THE SAME THEN:
30T=40(T-4)
30T=40T-160
30T-40T=160
-10T=-160
T=-160/-10
T=16 HOURS THEY WILL MEET.
PROOF:
30*16=40(16-4)
480=40*12
480=480